OpenStudy (anonymous):

1.) Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5). 2.) Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4). 3.) Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0). 4.) Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5). 5.) Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5).

OpenStudy (anonymous):

A) y = -2x - 1 B) y = -4x + 20 C) y = -5x + 15 D) y = -3x + 14 E) y = 2x

OpenStudy (texaschic101):

I will do a perpendicular and a parallel to show you how to do it... First you need to put the equations in y = mx + b form 1. x - 5y = -10 -5y = -x - 10 (multiply the equation by -1) 5y = x + 10 (now divide by 5) y = 1/5x + 2 Perpendicular lines have slopes that are the negative reciprocal. So the slope will be -5. y - y1 = m(x - x1) using slope -5 and points (2,5) y - 5 = -5(x - 2) y - 5 = -5x + 10 y = -5x + 10 + 5 y = -5x + 15 (answer C) 2. 10x - 5y = 8 -5y = -10x + 8 (multiply by -1) 5y = 10x - 8 (divide by 5) y = 2x -8/5 Parallel line have the same slope, so the slope will be 2. y - y1 = m(x - x1) using slope 2 and points (2,4) y - 4 = 2(x - 2) y - 4 = 2x - 4 y = 2x - 4 + 4 y = 2x (answer E) If you have any more problems with the rest of them, just tag me :)

OpenStudy (anonymous):

Thank you!!

OpenStudy (texaschic101):

do you understand enough to do the other problems ?

OpenStudy (anonymous):

Yes! Thanks!

OpenStudy (texaschic101):

glad I could help :)